# Host Galaxy Demographics Of Individually Detectable Supermassive Black-hole Binaries with Pulsar Timing Arrays

### v1

@@ 1 Introduction
There are a variety of interactions between the two massive black-holes (MBHs) and their ambient environment that can ‘harden’ it, i.e., dissipate orbital energy and lead to a reduction in black-hole separation. Broadly speaking, these mechanisms are stellar scattering, torques from a gaseous circumbinary disk, and the emission of gravitational waves [GWs; 11]; and at times three-body interactions from subsequent galaxy mergers [63, 13, 44, 17]. Once the binary separation is sufficiently small such that the mass enclosed is lower than the total mass of the black holes (typically at $\sim 1-10$ pc), the dynamics will be driven by the black holes themselves and they become gravitationally-bound as a “binary”. From the merging of host galaxies to formation of a binary may take several Gyrs [52], and significant uncertainties remain regarding the scattering of stars in the “final parsec” [67, 12, 58, 104].
At separations of $\lesssim 1$ pc, a circumbinary disk can form from accreting gas that is channeled into the central galactic regions as a result of the turbulent conditions following the galaxy merger [10]. If the black holes are of comparable mass, the torque exerted by the binary can carve out a cavity in the center of the circumbinary disk, leading to a depleted region within which the MBHB is surrounded by a dam of gas [75]. This dam can leak tendrils of gas onto the black holes, leading to the formation of circum-single disks (or ‘mini-disks’) around each [6, 41]. For decades it was believed that the torques exerted between the accretion disks and binary leads to energy and angular momentum extraction from the binary, further contributing to the system’s hardening [61, 40]. More recently, Miranda et al. [69] and Muñoz et al. [72] have shown that in some cases the presence of gas can expand the binary instead of contracting it. While this result has been confirmed by other groups [70, 31], its applicability to astrophysically realistic environments and particularly the high-masses of PTA detectable systems [e.g. 18] remains uncertain. Nonetheless, the presence of gas around the MBHs can give rise to several electromagnetic signatures, many producing variations related to the binary’s orbital period [e.g. 34, 26, 71, 25], these signatures could permit electromagnetic counterparts of any GW-detected system to be identified [59, 19, 27, 16, 24].
@@
There are three PTA collaborations that have been collecting observations for more than a decade: the European PTA [EPTA, 29], the North American Nanohertz Observatory for Gravitational waves [NANOGrav, 65], and the Parkes PTA [PPTA, 64]. Along with the recently-established Indian PTA [49], these four collaborations constitute the International Pulsar Timing Array [IPTA, 105], within which a coordinated effort to combine data and accelerate discoveries is being made [105, 76]. These collaborations are pursuing three main GW signal classes: $(i)$ a stochastic GW background created through the superposition of many SMBHB signals [84, 80, 110, 93], or of cosmological origin [e.g. 62, 20, 21]; $(ii)$ individually resolvable “continuous wave” SMBHB signals that may resound above the background [95, 88, 54]; and $(iii)$ GW burst signals that follow the ultimate merger of the binary [111, 22, 108, 102]. While the latter may be the only way for PTAs to observe the final coalescence of SMBHB evolution [3, 2], we do not consider this signal class here.
@@
Our main targets in this paper are resolvable GW signals from individual SMBHBs. While the stochastic GW background (SGWB) is now apparent from recent results [7, 1, 85, 32], PTAs with additional data are expected to be able to resolve several massive, nearby binaries out of the confusion background [89, 55, 9]. However, the sky-localization capability of PTAs for threshold binary detections is expected to be relatively poor ($\sim 10^{2}-10^{3}$ deg2) [91, 101, 37, 79], as is distance determination, rendering host-galaxy identification highly challenging [38, 79]. Our goal in this paper is to use cosmological hydrodynamical simulations combined with small-scale post-processing of binary dynamics and mock PTA configurations to forecast the demographics of galaxies that could host GW-detectable SMBHBs. By doing so, we aim to rank galaxies and thereby drastically reduce the number of plausible hosts that may lie within a GW localization volume from PTA binary detection. Statistical host identification will then trigger dedicated EM follow-up on the smaller subset of likely hosts to search for other morphological, spectroscopic, or photometric signatures of binarity.
@@
Several previous works have studied different aspects of this problem by conditioning predictions on real galaxy catalogs and simulations. Simon et al. [97] compiled a $90\%$-complete catalog of galaxies out to $\sim 200$ Mpc, injected putative binaries with properties based on galaxy-BH scaling relationships, then ranked potential hotspots based on the binary number density and directional GW power. Rosado & Sesana [87] built on this approach by adapting dark-matter halo clustering and merger metrics developed from the Millennium simulation to a value-added galaxy catalog based on SDSS DR7, of which the latter is limited in completeness and sky coverage. Mingarelli et al. [68] extended these approaches by populating massive galaxies with $D\lesssim 200$ Mpc from the 2MASS survey with putative binaries to assess likely hosts for near-future GW detections, and make empirical predictions for the level of statistical SWGB anisotropy from the local SMBHB population. [8] refined this approach by constructing a 2MASS-based catalog of $\sim 44,000$ galaxies out to $D\sim 500$ Mpc, and finding that hypothetical binaries within $216$ of those galaxies (mostly elliptical/early-type) lie within the GW-sensitivity volume of the NANOGrav $11$-year Dataset. Petrov et al. [79] injected putative SMBHBs into a subset of galaxies in different sky regions and at varying distances within this catalog, simulating the entire Bayesian detection, localization, and host-identification process, and applying subsequent cuts based on binary mass and distance consistency between the GW posterior and galaxy catalog values.
@@
From a simulation-driven perspective, Rosado et al. [90] generated many realizations of SMBHB populations using semi-analytic modeling, and assessed the probability of binary detection as a function of time and PTA configuration. They found that the SGWB was likely to be discovered prior to any individual binaries, but that the detection-probability weighted distributions of these individually-detectable systems favored $\mathcal{M}\sim 10^{9}-10^{10}M_{\odot}$, $z\sim 0.5$, and $f\lesssim 10$ nHz. Goldstein et al. [39] used a similar simulation scheme as Rosado et al. [90], and developed a statistical host-ranking metric based on null-stream GW detection statistics and galaxy-BH relationships to reduce the number of plausible hosts by several orders of magnitude. Kelley et al. [52] was the first PTA-study to employ the Illustris cosmological hydrodynamical simulation in which MBHs and galaxies co-evolve. Those results were used to predict properties of the GWB [53], and forecast the detection prospects and properties of individual MBHBs [56] in which they found only a weak dependence on astrophysical conditions such as binary eccentricity and environmental couplings. As expected, the PTA noise properties were found to have a much larger impact. Especially in the case of substantial pulsar red-noise, these results suggest that single binaries could be detected much sooner than previously expected (Kelley et al. 56; c.f. Rosado et al. 90).
@@
More recently, [28] used the Illustris simulation to probe whether the morphologies of recent mergers could indicate the presence of SMBHBs, finding that such evidence may persist for $\sim 500$ Myrs in PTA-type systems (defined as $\mathcal{M}>10^{8}M_{\odot}$). However, the lag time between galaxy merger and subsequent PTA-detection is such that EM follow-up is unlikely to detect these morphological disturbances. Additionally, they find that the specific star formation rate is enhanced in low-mass mergers, but mostly unaffected in high-mass mergers. However, these results are based on fixed post-merger lag times of 00, $0.5$, and $1$ Gyr, rather than taking into account distributions of coalescence timescales based on specific galaxy-BH interactions.
@@
This paper is laid out as follows. In Section 2 we outline the selection of SMBHBs and host galaxies from the Illustris simulation, as well the details of how we applied sub-grid binary dynamical evolution after the black holes were manually merged by the simulation upon reaching a gravitational softening separation ($\sim 1$ kpc). In Section 3 we describe the statistical framework for assessing the PTA detectability of each binary’s GW signal, and in Section 4 we characterize the demographics of galaxy observables for those galaxies that would be statistically-likely hosts for near-future PTA binary detections. We conclude and discuss future prospects in Section 5.
@@ 2.1 Merging supermassive black holes within Illustris
The Illustris simulations are known to accurately reproduce the distributions and properties of not only MBHs but also their host galaxies [96, 107]. In particular, Illustris accurately reproduces the range of galaxy types, and their appropriate correlations with brightness, color, star-formation rate, compactness, and environment [106, 36, 99]. Recently, the IllustrisTNG simulation suite was completed [74, 81, 100], which focuses on updated models for star-formation driven winds and AGN feedback aimed at better reproducing the inter- and circum- galactic media. The results from the IllustrisTNG simulations are still consistent with the original Illustris models for the characteristic scalar measurements of galaxies that we employ here (such as mass, magnitude, metallicity, etc). The TNG models increase the MBH ‘seed’ mass by an order of magnitude, drastically changing the distribution of MBH and MBH mergers properties, thus we have continued to use the original Illustris population to maintain continuity with previous GW studies [e.g. 14, 55, 57].
@@ 2.2 Evolving supermassive black-hole binaries into PTA sensitivity band
A detailed explanation of dynamical friction can be found in Antonini & Merritt [5], and the specifics of our implementation are found in Kelley et al. [52]. Stellar scattering is implemented following the formalism and fits to numerical scattering experiments from Sesana et al. [92] which assume a spherical and isotropic background of stars. When considering circumbinary disk torques, we adopt the three-zone thin accretion disk model from Haiman et al. [42] which is normalized to the accretion rates derived directly from Illustris [details in 52].
After evolving from the initial Illustris separation due to these environmental factors the binaries may reach a separation that is small enough such that the hardening will be dominated by the emission of GWs. The GW hardening equations may be found in Peters [78], and the details of the post processing follows that of Kelley et al. [52].
@@ 2.3 Tracking host galaxies over cosmic time within Illustris
In order to access the properties of a specific galaxy in the simulation, one needs the subhalo ID and the snapshot number (in our case a subhalo in Illustris is treated the same as a galaxy). We then track the evolution of the galaxy over time using galaxy merger trees [‘sublink’; 86]. Having the subhalo ID and snapshot directly after the black holes “merge” in Illustris allows us to follow the merger tree into the future of the subhalo to the age of the Universe at which it would produce GWs in the PTA sensitivity band. We choose the first simulation snapshot following this time to analyze the galaxy properties. This gives the subhalo ID and snapshot of the galaxy in Illustris when our post-processing deems it is detectable.
To find the right subhalo, we had to navigate the subhalo merger trees. The sublink merger tree for the main descendant branch (sublink_mdb) includes arrays of all the snapshot numbers of the descendants of this system, carrying with each of them the ID number at that snapshot. Figure 1 gives a visual depiction of this idea. The index of the correct system in the array of snapshots is the system’s subhalo ID for that snapshot. There are four systems where the merger tree skips over the snapshot where the desired system should become PTA detectable. This is a workaround in Illustris to address situations in which one subhalo is passing through another at the snapshot of interest [86]. These are removed from the set of systems we study, but should have a negligible impact on our analysis.
Figure 1: The dark blue circled system is the subhalo immediately preceeding the MBH ‘merger’ event in Illustris (black line), for which we are retrieving the merger tree. We are interested in determining the descendant galaxy (green circle) at the time of PTA detectability (grey line). The subhalos within the green box are along the intial subhalo’s descendent branch.
The properties of the host galaxy that we examine are the redshift, galaxy mass, stellar metallicity, specific star formation rate, and color–magnitude relationship. Additionally, we study the chirp mass and frequency of the binary to compare with the other properties of the host galaxy. In order to understand the properties of the host galaxies of systems that are the most likely to be detected, the property histograms must be weighted by some metric to indicate their likelihood of being detected. We choose the detection probability [DP; 90] to weight our system, but we compare the DP weighting to weighting by the strain, signal-to-noise, and simply a frequency cutoff for MBH systems which assumes that all systems in the PTA band are detectable.
Figure 2: Gravitational-wave characteristic strain spectra from Monte Carlo realizations of SMBHB populations resampled from Illustris. Five realizations are shown in the left panel, with each having a different color. The dashed gray line is the expected power-law GWB spectrum assuming binary evolution solely by GW emission, such that $h_{c}\propto f^{-2/3}$. In the right panel only a single realization is shown, but with the strain spectrum dissected to show the total spectrum as well as the remnant after removing the loudest binary from each frequency bin.

### v2

@@ 1 Introduction
There are a variety of interactions between the two massive black-holes (MBHs) and their ambient environment that can ‘harden’ it, i.e., dissipate orbital energy and lead to a reduction in black-hole separation. Broadly speaking, these mechanisms are stellar scattering, torques from a gaseous circumbinary disk, and the emission of gravitational waves [GWs; 14]; and at times three-body interactions from subsequent galaxy mergers [74, 17, 50, 21]. Once the binary separation is sufficiently small such that the mass enclosed is lower than the total mass of the black holes (typically at $\sim 1-10$ pc), the dynamics will be driven by the black holes themselves and they become gravitationally-bound as a “binary”. From the merging of host galaxies to formation of a binary may take several Gyrs [61], and significant uncertainties remain regarding the scattering of stars in the “final parsec” [78, 16, 69, 119].
At separations of $\lesssim 1$ pc, a circumbinary disk can form from accreting gas that is channeled into the central galactic regions as a result of the turbulent conditions following the galaxy merger [13]. If the black holes are of comparable mass, the torque exerted by the binary can carve out a cavity in the center of the circumbinary disk, leading to a depleted region within which the MBHB is surrounded by a dam of gas [86]. This dam can leak tendrils of gas onto the black holes, leading to the formation of circum-single disks (or ‘mini-disks’) around each [7, 47]. For decades it was believed that the torques exerted between the accretion disks and binary leads to energy and angular momentum extraction from the binary, further contributing to the system’s hardening [72, 46]. More recently, Miranda et al. [80] and Muñoz et al. [83] have shown that in some cases the presence of gas can expand the binary instead of contracting it. While this result has been confirmed by other groups [81, 37, 111], its applicability to astrophysically realistic environments and particularly the high-masses of PTA detectable systems [e.g. 22, 112] remains uncertain. Nonetheless, the presence of gas around the MBHs can give rise to several electromagnetic signatures, many producing variations related to the binary’s orbital period [e.g. 40, 31, 82, 29], these signatures could permit electromagnetic counterparts of any GW-detected system to be identified [70, 23, 33, 20, 28].
@@
There are three PTA collaborations that have been collecting observations for more than a decade: the European PTA [EPTA, 35], the North American Nanohertz Observatory for Gravitational waves [NANOGrav, 76], and the Parkes PTA [PPTA, 75]. Along with the recently-established Indian PTA [57], these four collaborations constitute the International Pulsar Timing Array [IPTA, 120], within which a coordinated effort to combine data and accelerate discoveries is being made [120, 88]. These collaborations are pursuing three main GW signal classes: $(i)$ a stochastic GW background created through the superposition of many MBHB signals [96, 92, 126, 106], or of cosmological origin [e.g. 73, 24, 25]; $(ii)$ individually resolvable “continuous wave” MBHB signals that may resound above the background [108, 100, 64]; and $(iii)$ GW burst signals that follow the ultimate merger of the binary [127, 26, 124, 117]. While the latter may be the only way for PTAs to observe the final coalescence of MBHB evolution [3, 2], we do not consider this signal class here.
@@
Our main targets in this paper are resolvable GW signals from individual MBHBs. While the stochastic GW background (SGWB) is now apparent from recent results [8, 1, 97, 38], PTAs with additional data are expected to be able to resolve several massive, nearby binaries out of the confusion background [101, 65, 10]. However, the sky-localization capability of PTAs for threshold binary detections is expected to be relatively poor ($\sim 10^{2}-10^{3}$ deg2) [104, 116, 43, 91], as is distance determination, rendering host-galaxy identification highly challenging [44, 91]. Our goal in this paper is to use cosmological hydrodynamical simulations combined with small-scale post-processing of binary dynamics and mock PTA configurations to forecast the demographics of galaxies that could host GW-detectable MBHBs. By doing so, we aim to rank galaxies and thereby drastically reduce the number of plausible hosts that may lie within a GW localization volume from PTA binary detection. Statistical host identification will then trigger dedicated EM follow-up on the smaller subset of likely hosts to search for other morphological, spectroscopic, or photometric signatures of binarity.
@@
Several previous works have studied different aspects of this problem by conditioning predictions on real galaxy catalogs and simulations. Simon et al. [110] compiled a $90\%$-complete catalog of galaxies out to $\sim 200$ Mpc, injected putative binaries with properties based on galaxy-BH scaling relationships, then ranked potential hotspots based on the binary number density and directional GW power. Rosado & Sesana [99] built on this approach by adapting dark-matter halo clustering and merger metrics developed from the Millennium simulation to a value-added galaxy catalog based on SDSS DR7, of which the latter is limited in completeness and sky coverage. Mingarelli et al. [79] extended these approaches by populating massive galaxies with $D\lesssim 200$ Mpc from the 2MASS survey with putative binaries to assess likely hosts for near-future GW detections, and make empirical predictions for the level of statistical SWGB anisotropy from the local MBHB population. [9] refined this approach by constructing a 2MASS-based catalog of $\sim 44,000$ galaxies out to $D\sim 500$ Mpc, and finding that hypothetical binaries within $216$ of those galaxies (mostly elliptical/early-type) lie within the GW-sensitivity volume of the NANOGrav $11$-year Dataset. Petrov et al. [91] injected putative MBHBs into a subset of galaxies in different sky regions and at varying distances within this catalog, simulating the entire Bayesian detection, localization, and host-identification process, and applying subsequent cuts based on binary mass and distance consistency between the GW posterior and galaxy catalog values.
@@
From a simulation-driven perspective, Rosado et al. [102] generated many realizations of MBHB populations using semi-analytic modeling, and assessed the probability of binary detection as a function of time and PTA configuration. They found that the SGWB was likely to be discovered prior to any individual binaries, but that the detection-probability weighted distributions of these individually-detectable systems favored $\mathcal{M}\sim 10^{9}-10^{10}M_{\odot}$, $z\sim 0.5$, and $f\lesssim 10$ nHz. Goldstein et al. [45] used a similar simulation scheme as Rosado et al. [102], and developed a statistical host-ranking metric based on null-stream GW detection statistics and galaxy-BH relationships to reduce the number of plausible hosts by several orders of magnitude. Kelley et al. [61] was the first PTA-study to employ the Illustris cosmological hydrodynamical simulation in which MBHs and galaxies co-evolve. Those results were used to predict properties of the GWB [62], and forecast the detection prospects and properties of individual MBHBs [66] in which they found only a weak dependence on astrophysical conditions such as binary eccentricity and environmental couplings. As expected, the PTA noise properties were found to have a much larger impact. Especially in the case of substantial pulsar red-noise, these results suggest that single binaries could be detected much sooner than previously expected (Kelley et al. 66; c.f. Rosado et al. 102).
@@
More recently, [34] used the Illustris simulation to probe whether the morphologies of recent mergers could indicate the presence of MBHBs, finding that such evidence may persist for $\sim 500$ Myrs in PTA-type systems (defined as $\mathcal{M}>10^{8}M_{\odot}$). However, the lag time between galaxy merger and subsequent PTA-detection is such that EM follow-up is unlikely to detect these morphological disturbances. Additionally, they find that the specific star formation rate is enhanced in low-mass mergers, but mostly unaffected in high-mass mergers. However, these results are based on fixed post-merger lag times of 00, $0.5$, and $1$ Gyr, rather than taking into account distributions of coalescence timescales based on specific galaxy-BH interactions.
While galaxy morphologies may be the most robust indicators of recent galaxy mergers, they are only feasible for a relatively small number of nearby galaxies. Due to the large distances and thus numbers of possible host galaxies within typical PTA sensitive volumes, it will be crucial to leverage deep, all-sky photometric surveys such as the upcoming Vera Rubin Observatory LSST [54]. The bulk of the billions of galaxies cataloged by LSST will only have point-like measurements of colors and brightness. Making selections in the standard, galaxy color-magnitude space  [^1] The galaxy color-magnitude diagram is typically presented as r-band absolute magnitude (a proxy for total mass) plotted against g-band minus r-band color (‘g-r‘, a proxy for typical stellar ages). [e.g. 15] is thus the most valuable initial cut on possible host galaxies.
@@
This paper is laid out as follows. In Section 2 we outline the selection of MBHBs and host galaxies from the Illustris simulation, as well the details of how we applied sub-grid binary dynamical evolution after the black holes were manually merged by the simulation upon reaching a gravitational softening separation ($\sim 1$ kpc). In Section 3 we describe the statistical framework for assessing the PTA detectability of each binary’s GW signal, and in Section 4 we characterize the demographics of galaxy observables for those galaxies that would be statistically-likely hosts for near-future PTA binary detections. We conclude and discuss future prospects in Section 5.
@@ 2.1 Merging supermassive black holes within Illustris
The Illustris simulations are known to accurately reproduce the distributions and properties of not only MBHs but also their host galaxies [109, 122]. In particular, Illustris accurately reproduces the range of galaxy types, and their appropriate correlations with brightness, color, star-formation rate, compactness, and environment [121, 42, 114]. Recently, the IllustrisTNG simulation suite was completed [85, 93, 115], which focuses on updated models for star-formation driven winds and AGN feedback aimed at better reproducing the inter- and circum- galactic media. The results from the IllustrisTNG simulations are still consistent with the original Illustris models for the characteristic scalar measurements of galaxies that we employ here (such as mass, magnitude, metallicity, etc). The TNG models increase the MBH ‘seed’ mass by an order of magnitude, drastically changing the distribution of MBH and MBH mergers properties, thus we have continued to use the original Illustris population to maintain continuity with previous GW studies [e.g. 18, 65, 67].
@@ 2.2 Evolving supermassive black-hole binaries into PTA sensitivity band
A detailed description of dynamical friction can be found in Antonini & Merritt [6], and the specifics of our implementation are found in Kelley et al. [61]. Stellar scattering is implemented following the formalism and fits to numerical scattering experiments from Sesana et al. [105] which assume a spherical and isotropic background of stars. When considering circumbinary disk torques, we adopt the three-zone thin accretion disk model from Haiman et al. [48] which is normalized to the accretion rates derived directly from Illustris. After evolving due to these environmental factors, the binaries may reach a separation that is small enough such that the hardening will be dominated by the emission of GWs. The GW hardening equations may be found in Peters [90].
Binary evolution en masse typically takes billions of years, with only a fraction of binary systems coalescing before redshift zero. Total binary lifetimes depend sensitively on not only on the masses of the two MBHs, but also the structure and composition of the post-merger host galaxy. In particular, there are two standard bottlenecks for mergers where some MBHBs can ‘stall’. The first is at kiloparsec-scale separations, where tidal stripping of the secondary galaxy can substantially slow dynamical friction, and lead to the secondary MBH being deposited in the primary galaxy’s outskirts. This typically occurs in unequal mass-ratio mergers [$q\equiv m_{2}/m_{1}\lesssim 0.1$ 61]. The second is at parsec separations, the so-called ‘final-parsec problem’ [78], where the population of stars able to scatter with (and extract energy from) the binary becomes depleted. There is now a general consensus that most galaxies are able to sufficiently replenish this population of stars [e.g. 69, 119], although MBHBs in low total-mass galaxies and those with shallow inner stellar density profiles can still stall [61]. Accurate models for binary evolution thus encode important correlations between MBHs/galaxies and which binaries are able to reach the small separations corresponding to detectable GW emission.
Our evolution models do not account for ‘triple’ MBH interactions [17, 51], although these are expected to be subdominant for MBHB mergers overall [103]. Additionally, while we do account for circumbinary disk torques, we do not include MBH growth during inspiral [111]; this should not significantly effect binary lifetimes in the PTA regime [22, 112].
@@ 2.3 Tracking host galaxies over cosmic time within Illustris
After the two MBH particles are ‘merged’ in Illustris, we identify the host-galaxy in the subsequent simulation snapshot, and the post-processing binary evolution model (described above) yields the age of the Universe at which each binary reaches the PTA sensitivity band. We then track the evolution of the galaxy using merger trees [‘sublink’; 98], until we reach the snapshot immediately following the PTA-detectable epoch. The sublink merger tree for the main descendant branch (sublink_mdb) includes arrays of all the snapshot numbers of the descendants of this system, carrying with each of them the ID number at that snapshot. The index of the correct system in the array of snapshots is the system’s subhalo ID for that snapshot. There are four systems where the merger tree skips over the snapshot where the desired system should become PTA detectable. This is a workaround in Illustris to address situations in which one subhalo is passing through another at the snapshot of interest [98]. These are removed from the set of systems we study, but should have a negligible impact on our analysis.
The properties of the host galaxy that we examine are the redshift, galaxy mass, stellar metallicity, specific star formation rate, and color–magnitude relationship. Additionally, we study the chirp mass and frequency of the binary to compare with the other properties of the host galaxy. In order to understand the properties of the host galaxies of systems that are the most likely to be detected, the property histograms must be weighted by some metric to indicate their likelihood of being detected. We choose the detection probability [DP; 102] to weight our system, but we compare the DP weighting to weighting by the strain, signal-to-noise, and simply a frequency cutoff for MBH systems which assumes that all systems in the PTA band are detectable.
Figure 1: Gravitational-wave characteristic strain spectra from Monte Carlo realizations of MBHB populations resampled from Illustris. Five realizations are shown in the left panel, with each having a different color. The dashed gray line is the expected power-law GWB spectrum assuming binary evolution solely by GW emission, such that $h_{c}\propto f^{-2/3}$. In the right panel only a single realization is shown, but with the strain spectrum dissected to show the total spectrum as well as the remnant after removing the loudest binary from each frequency bin.